OpenStudy (anonymous):

What is the solution to the equation? 32 - 6x + 7x = -15 + 28 x=________

OpenStudy (anonymous):

so do this -15 + 28 what does it equal

OpenStudy (anonymous):

Simplifying 32 + -6x + 7x = -15 + 28 Combine like terms: -6x + 7x = 1x 32 + 1x = -15 + 28 Combine like terms: -15 + 28 = 13 32 + 1x = 13 Solving 32 + 1x = 13 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-32' to each side of the equation. 32 + -32 + 1x = 13 + -32 Combine like terms: 32 + -32 = 0 0 + 1x = 13 + -32 1x = 13 + -32 Combine like terms: 13 + -32 = -19 1x = -19 Divide each side by '1'. x = -19 Simplifying x = -19